OpenStudy (anonymous):

Inverse Laplace Transform (2s^2+4s+8)/(s^2+4)(s^2+2s+1)

OpenStudy (anonymous):

@AccessDenied can u tall me the way how to do this equation ?

OpenStudy (zarkon):

partial fraction...then find inverse

OpenStudy (anonymous):

can u show me it plz

OpenStudy (anonymous):

Let\[\frac{As+B}{s^2+4}+\frac{Cs+D}{s^2+2s+1}=\frac{2s^2+4s+8}{(s^2+4)(s^2+2s+1)}\] So, we get \[(As+B)(s^2+2s+1)+(Cs+D)(s^2+4) \equiv 2s^2+4s+8 \] \[(A+C)s^3+(2A+B+D)s^2+(A+2B+4C)s+(B+4D) \equiv 2s^2+4s+8 \]Hence, we have four equations in four unknown. They are: A+C =0 2A+B+D = 2 A+2B+4C = 4 B+4D = 8 Solve A,B,C,D, then you'll be done with the partial fractions part.